### 03 Feb 2024 v1

# Unveiling Latent Causal Rules: A Temporal Point Process Approach for Abnormal Event Explanation

Abstract
In high-stakes systems such as healthcare, it is critical to understand the causal reasons behind unusual events, such as sudden changes in patient’s health. Unveiling the causal reasons helps with quick diagnoses and precise treatment planning. In this paper, we propose an automated method for uncovering “if-then” logic rules to explain observational events. We introduce temporal point processes to model the events of interest, and discover the set of latent rules to explain the occurrence of events. To achieve this, we employ an Expectation-Maximization (EM) algorithm. In the E-step, we calculate the likelihood of each event being explained by each discovered rule. In the M-step, we update both the rule set and model parameters to enhance the likelihood function’s lower bound. Notably, we optimize the rule set in a differential manner. Our approach demonstrates accurate performance in both discovering rules and identifying root causes. We showcase its promising results using synthetic and real healthcare datasets.
1 Introduction
Detecting and understanding abnormal events is crucial in various fields. In healthcare, quickly spotting and diagnosing sudden patient deterioration can save lives. In e-commerce, detecting odd user behavior may stop fraud and ensure safe shopping experiences. In manufacturing, finding irregularities in production processes can make operations more efficient and reduce downtime. These examples highlight the importance of abnormal event detection for enhancing safety, security, and efficiency in various areas.
However, effectively implementing abnormal event detection comes with challenges. A major obstacle is the lack of labeled abnormal event data during training. Labeling events as normal or abnormal is often expensive and labor-intensive. Consider the context of credit card transactions. When the system encounters an unusually large transaction, determining its underlying rationale can be intricate. It may indeed be a legitimate and routine transaction, or it could be indicative of fraudulent activity. Identifyineng the true reasons require thorough investigation, which is both time and resource consuming. In our paper, we address this issue by introducing a rule-based probabilistic approach for predicting and explaining events. We treat event’s explanatory reasons as hidden variables and use an EM algorithm to learn rules that explain events without explicit label information regarding anomaly. For each observed event, the triggered rules will provide insight into the most probable causal factors and determine whether the event should be flagged as abnormal or normal.
We introduce temporal point process to model events of interest in continuous time. The time-to-event is treated as a random variable characterized by the intensity function, indicating the occurrence rate of the event in real time. We formulate the intensity function as a mixture model, where each component is defined by an “if-then” rule factor. We assume that the “if” condition should occur before the “then” part, and the “if” condition also specifies the temporal order constraints that the relevant logic variables must meet, such as “if $X_{1}$ and $X_{2}$ are true and $X_{1}$ happens before $X_{2}$, then $Y$ is true.” These explanatory temporal logic rules will be discovered from data, which compete to influence when an event occurs. When an event happens, the specific rule that is triggered based on preceding events will provide an explanation for its occurrence.
As mentioned above, we formulate an EM algorithm to simultaneously learn the rule set and temporal point processes model parameters. In the E-step, we calculate the assignment probability of each event to its causal rules based on our current estimations of the model parameters and the rule set. In the M-step, we update the rule set and model parameters to maximize the expected likelihood calculated in E-step.
Notably, learning the rule set poses a formidable challenge due to its inherently combinatorial nature. It requires selecting the proper subsets of logic variables and then considering their temporal order relations to create a logic rule. This variable selection process is both time-consuming and non-differentiable. In our paper, we propose to solve a relaxed continuous optimization problem by directly modeling the logic variable selection probability in forming a rule. Specifically, each rule is generated by a reparameterized K-subset sampling with the Gumbel noise injected to make the rule sampling process differentiable [18]. We further use a padding trick to make the generated rules with various lengths not exceeding K. In the M-step, both the rule set and the model parameters can be learned end-to-end in a differentiable way.
Our contributions can be summarized as follows.
(i) We leverage the continuous-time temporal point process modeling framework and formulate the intensity function as a rule-informed mixture model. Our rule-based probabilistic model is inherently interpretable, ensuring safe and reliable event prediction and explanation.
(ii) We design an EM learning algorithm that can accommodate the latent logical reasons for each event. Our overall learning algorithm is efficient and differentiable, which can automatically discover the set of explanatory logic rules from the population data and determine the most likely logical reason for each event.
(iii) We conduct empirical evaluations of our algorithm using both synthetic and real datasets, demonstrating its strong performance in rule discovery and root cause determination.
2 Related Work
We will compare our method with some existing works from the following aspects.
Temporal Point Process (TPP) Models
TPP models provide an elegant tool for modeling the time intervals or time-to-event in continuous time. Current research in this area primarily focuses on enhancing the flexibility of intensity functions. For example, [4] introduced the RMTPP, a neural point process model that employs a Recurrent Neural Network to model the intensity function. [12] further improved RMTPP by developing a continuous-time RNN. [23] and [21] leveraged the self-attention mechanism to capture the long-term dependencies of events. However, relying on flexible black-box models, which can excel at event prediction with sufficient training data, presents challenges. These models lack interpretability, making it difficult to understand the occurrence of events. In scenarios like abnormal event detection, the objective goes beyond merely predicting unusual events; it also involves provding causal explanations for these occurrences. The absence of interpretability renders black-box models unsuitable for high-stakes systems where transparency is critical.
The contemporary AI perspective underscores the value of crafting intrinsically interpretable models over relying solely on explaining black-box models [14]. Along this line of research, recently, [11; 10] proposed to design a rule-informed intensity function, trying to add the interoperability. However, their approaches lack the capability to provide detailed logical explanations for individual events, in contrast to our method. Additionally, their rule learning process is non-differentiable, relying on a branch and bound algorithm.
Rule Set Mining
Rule set discovery in an unsupervised fashion has long been a classic topic. Traditional approaches include Itemset Mining Methods like Apriori [1] and NEclatcloesed [2], which target frequent itemset identification without accounting for event order. Sequential Pattern Mining methods, typified by CM-SPADE [6] and VGEN [7], aim to unveil temporal relationships but they lack precision and cannot handle timestamp information. In addition to the unsupervised methods, Inductive Logic Programming (ILP) [16] provides a powerful supervised rule learning method, yet requiring balanced positive and negative examples to have a good performance. In contrast, our method leverages fine-grained temporal information to learn temporal logic rules from data and in an unsupervised manner.
3 Problem Setup
We define a set of logic variables (or predicates) as $\mathcal{X}$, where each variable $X_{u}\in\mathcal{X}$ is a boolean variable. Our target event, such as a sudden change in patient’s health or an unusual large transaction, is denoted as $Y$ and $Y\in\mathcal{X}$. We add a temporal dimension to the logic variables and each grounded logic variables becomes a sequence of spiked events, denoted as $\{X_{u}(t)\}_{t\geq 0}$, where each $X_{u}(t)\in\{0,1\}$, $\forall t$. Specifically, $X_{u}(t)$ jumps to 1 (i.e., true) at the time step when the event happens. In our problem, each data sample is a $|\mathcal{X}|$-dimensional multivariate temporal processes, denoted as $\mathcal{H}_{t}=\{X_{u}(t)\}_{u=1,\dots,|\mathcal{X}|}$.
We are interested in modeling and providing insight into the occurrence of $\{Y(t)\}_{t\geq 0}$. We treat the time intervals as random variables and the duration until event $Y$ happens is characterized by the conditional intensity function, denoted as $\lambda(t\mid\mathcal{H}_{t-})$ where $t-$ means up to $t$ but not include $t$. By definition, $\lambda(t\mid\mathcal{H}_{t-})dt=\mathbb{E}[N([t,t+dt])\mid\mathcal{H}_{t-}]$, where $N(([t,t+dt])$ denotes the number of events occurring in the interval $[t,t+dt)$. Given the occurrence time of event $Y$, i.e., $(t_{1},\dots,t_{n})$, the joint likelihood function is given by
$\quad p(t_{1},\dots,t_{n})=\prod_{i=1}^{n}p^{*}(t_{i}),\quad\text{with}$
[…]
Here, to simplify the notation, we denote $p^{*}(t):=p(t\mid\mathcal{H}_{{t}-})$ and $\lambda^{*}(t):=\lambda(t\mid\mathcal{H}_{t-})$. In this paper, we will model $\lambda^{*}(t)$ using the rule-based features, through which we aim to understand the underlying logical reasons to explain the occurrence of target events. Specifically, our goals include:
(i) Discover the rule set :=$\mathcal{F}:=\{f_{1},f_{2},...,f_{H}\}$ from the entire events $\mathcal{H}_{t_{n}}$ up to $t_{n}$, with each rule having a general form
$f:\quad Y\leftarrow\left(\bigwedge\limits_{X_{u}\in\mathcal{X}_{f}}X_{u}\right)\bigwedge\left(\bigwedge\limits_{X_{u},X_{v}\in\mathcal{X}_{f}}r_{j}(X_{u},X_{v})\right)$ (2)
where $\mathcal{X}_{f}$ is the set of body predicates of rule $f$. ${r_{j}(X_{u},X_{v})}$ is a set of temporal relations between pairs of predicates and can take any of the three types, i.e., “Before”, “Equal”, and “After”, and can also be none, meaning there are no temporal order constants existing for $X_{u}$ and $X_{v}$.
(ii) Infer the assignment of each occurred target event $\{Y(t_{1}),\dots,Y(t_{n})\}$ to the following two groups:
  • spontaneous events (i.e., cannot be explained by any logic rule or it’s meaningless to be explained);
  • can be explained by one logic rule from the discovered rule set =$\mathcal{F}=\{f_{1},f_{2},...,f_{H}\}$.
4 Model
We first describe how to construct the intensity function for $Y$ based on the logic-informed features. For a general rule as defined in Eq. (2), we can construct the rule-informed feature as
$\phi_{f}(\mathcal{H}_{t})=\prod_{X_{u}\in\mathcal{X}_{f}}X_{u}(t_{u})\cdot\prod_{X_{u},X_{v}\in\mathcal{X}_{f}}r_{j}(X_{u},X_{v})\in\{0,1\}$ (3)
where we need to check from $\mathcal{H}_{t}$ whether each $X_{u}\in\mathcal{X}_{f}$ has been once grounded as true and whether the temporal relation grounded by their times holds (for multiple events we choose the last one). The temporal relation $r_{j}(X_{u},X_{v})$ is grounded according to its type,“Before”, “Equal”, or “After” (denoted as $R_{b},R_{e},R_{a}$ respectively) by the event times as follows,
$R_{b}\left(t_{u},t_{v}\right)$ $=\mathbbm{1}\left\{t_{u}-t_{v}<-\delta\right\}$
$R_{e}\left(t_{u},t_{v}\right)$ $=\mathbbm{1}\left\{\left|t_{u}-t_{v}\right|<\delta\right\}$
$R_{a}\left(t_{u},t_{v}\right)$ $=\mathbbm{1}\left\{t_{u}-t_{v}>\delta\right\}$
where $\delta$ (can be 0) is specified as the time tolerance. If the body condition of $f$ is grounded to be true given history, then the feature $\phi_{f}(\mathcal{H}_{t})$ is 1, otherwise it is 0. A better illustration can be found in Fig. 1.
Figure 1: Illustration of the feature construction. For the logic rule: $f:Y\leftarrow X_{1}\bigwedge X_{2}\bigwedge X_{3}$, whenever the body condition becomes trues, the rule gets fired, as a result the feature $\phi_{f}$ becomes 1, and the intensity function of the head predicate is boosted by $r_{f}$.
Given the rule-informed feature defined in Eq. (3), we can build our intensity model for $\{Y(t)\}_{t\geq 0}$ as follows. For the $i$-th event in the dataset, by considering all the rules =$\mathcal{F}=\{f_{1},f_{2},...,f_{H}\}$, we model the intensity function as a mixture-of-components form,
$\lambda^{*}(t_{i}\mid\bm{z}_{i})=b_{0}+\sum_{h=1,\dots,H}\gamma_{h}z_{ih}\phi_{f_{h}}(\mathcal{H}_{t_{i}})$ (4)
where $b_{0}$ is the base term, $h$ is the index of the rule, and $\gamma_{h}>0$ denotes the impact weight of each rule $f_{h}$. The latent variable $\bm{z}_{i}=[z_{ih}]_{h=1,\dots,H}$ is a one-hot vector. When $z_{ih}=1$, it means event $i$ is due to the $h$-th rule factor, which provides logical reasons to explain its occurrence. We assume the categorical $\bm{z}_{i}$ has a distribution $\bm{\pi}=[\pi_{h}]$, which indicates the probability of a rule appearing in the population and is our learnable parameter.
In the next section, we design an EM algorithm, which aims to jointly learn the rule set and point process model parameters from data, and simultaneously perform the inference of each $\bm{z}_{i}$.
5 Learning: EM Algorithm
To accommodate the latent rule assignment $\bm{z}_{i}$ for each event, it is natural to use an EM learning algorithm.
  • (i)
    In the E-step, the posterior probability over $z_{i}=[z_{ih}]_{h=0,\dots,H}$ for each event $Y^{i}$ is computed. We use $z_{i0}$ to indicate that the event cannot be explained by any discovered rule.
  • (ii)
    In the M-step, both the rule set and the point process model parameters are updated by maximizing the expected likelihood found in E-step. Denote all the learnable model parameters as
    $\theta=[b_{0},[\gamma_{h}]_{h\in[1,\dots,H]},\bm{\pi},\mathcal{F}].$
Notably, we devise a series of techniques in EM to enhance the stability of the optimization process. The E-step has been successfully resolved in a closed-form solution. The M-step is more involved, for which we employ three-fold approximations and reparameterization to facilitate the end-to-end differentiable training.
Complete Data Likelihood
To simplify the derivation, let’s start with the $i$-th event $Y(t_{i})$ and consider its complete data likelihood. The following derivations can be easily extended to the entire event sequence.
For an occurred event $Y(t_{i})$ at time $t_{i}$, given the rule index $\bm{z}_{i}=[z_{ih}]_{h=0,\dots,H}$, we can write the likelihood for the complete data $\{(t_{i},\bm{z}_{i})\}$ as
$p^{*}_{\theta}(t_{i},\bm{z}_{i})=p^{*}_{\theta}(t_{i}\mid\bm{z}_{i})p(\bm{z}_{i}),\quad\text{where}$
$p(z_{ih}=1)=\pi_{h},\quad\text{with}\quad 0\leq\pi_{h}\leq 1,\,\sum_{h=0}^{H}\pi_{h}=1$
$p^{*}_{\theta}(t_{i}\mid z_{ih}=1)$
[…]
The complete data log-likelihood $\log p^{*}_{\theta}(t_{i},\bm{z}_{i})$ is easier to optimize compared to the log-likelihood of data, computed as
$\ell(\theta)=\log\left[\sum_{\bm{z}_{i}}p_{\theta}^{*}(t_{i},\bm{z}_{i})\right]$ (6)
which requires the marginalization of the latent variable $\bm{z}_{i}$.
5.1 E-step: Compute Posterior $Q^{new}(\bm{z}_{i})$
To derive the lower bound of the log-likelihood, let’s denote the a distribution for $\bm{z}_{i}=[z_{ih}]_{h=1,\dots,H}$ as $Q(\bm{z}_{i})$ and $Q\left(\bm{z}_{i}\right)\geq 0$ and $\sum_{\bm{z}_{i}}Q\left(\bm{z}_{i}\right)=1$. The above constraints make $Q(\bm{z}_{i})$ a valid probability mass function.
Now we can create a lower bound for log-likelihood function
$\ell(\theta)=\log\mathbb{E}_{Q(\bm{z}_{i})}\left[\frac{p^{*}_{\theta}(t_{i},\bm{z}_{i})}{Q(\bm{z}_{i})}\right]\geq\mathbb{E}_{Q(\bm{z}_{i})}\left[\log\frac{p_{\theta}^{*}\left(t_{i},\bm{z}_{i}\right)}{Q\left(\bm{z}_{i}\right)}\right]$
where the inequality is due to the Jensen’s inequality.
In the current iteration of the EM algorithm, suppose we aim to find $Q^{new}(\bm{z}_{i})$, so that the lower bound equals to log-likelihood $\ell\left(\theta^{old}\right)$. According to Jensen’s inequality, the equality between the lower bound and $\ell\left(\theta^{old}\right)$ holds if $\frac{p_{\theta^{old}}^{*}\left(t_{i},\bm{z}_{i}\right)}{Q^{new}\left(\bm{z}_{i}\right)}$ is a constant. In other words, $Q^{new}(\bm{z}_{i})$ needs to take the posterior of $\bm{z}_{i}$ to make the lower bound tight, i.e.,
E-step: $Q^{new}\left(z_{ih}=1\right)=\frac{p^{*}_{\theta^{old}}\left(t_{i},z_{ih}=1\right)}{\sum_{h^{\prime}=0}^{H}p^{*}_{\theta^{old}}\left(t_{i},z_{ih^{\prime}}=1\right)}$
$=\frac{\pi_{h}^{old}p^{*}_{\theta^{old}}(t_{i}\mid z_{ih}=1)}{\sum_{h^{\prime}=0}^{H}\pi^{old}_{h^{\prime}}p^{*}_{\theta^{old}}(t_{i}\mid z_{ih^{\prime}}=1)}$ (7)
with a more detailed calculation can be found in Eq. (5).
5.2 M-step: Optimize $\theta^{\text{new}}$
For the M-step, we aim to update the model parameters $\theta$ (including the rule set $\mathcal{F}$) to maximize the expected log-likelihood computed in the E-step. which is a lower bound of the log-likelihood, i.e.,
$\theta^{new}:=\arg\max_{\theta}\underbrace{\mathbb{E}_{Q^{new}(\bm{z}_{i})}\left[\log\frac{p_{\theta}^{*}\left(t_{i},\bm{z}_{i}\right)}{Q^{new}\left(\bm{z}_{i}\right)}\right]}_{\text{``energy"}}$
which is equal to solve (since $Q^{new}(\bm{z}_{i})$ does not depend on $\theta$)
$\theta^{new}:=\arg\max_{\theta}\mathbb{E}_{Q^{new}(\bm{z}_{i})}\left[\log p_{\theta}^{*}\left(t_{i},\bm{z}_{i}\right)\right]$ (8)
where $\log p_{\theta}^{*}\left(t_{i},\bm{z}_{i}\right)$ is the complete data log-likelihood. We further partition $\theta$ as $\theta:=[\theta_{0},\mathcal{F}]$ where $\theta_{0}$ contains the model parameters excluding $\mathcal{F}$. Then we rewrite the above formulation as
$\max_{\theta_{0}}\max_{\mathcal{F}}\mathbb{E}_{Q^{new}(\bm{z}_{i})}\left[\log p^{*}_{\theta_{0},\mathcal{F}}\left(t_{i},\bm{z}_{i}\right)\right]$ (9)
and we will alternate the optimization of $\theta_{0}$ and $\mathcal{F}$.
5.2.1 Optimization wrt $\mathcal{F}$
We first discuss how to update $\mathcal{F}$ given current $\theta_{0}.$ We aim to propose a differentiable method to learn the rule set. To achieve this goal, we will employ three-fold approximations or reparameterization:
  • (i)
    Continuous Relaxation of the Rule Learning: We formulate the rule learning problem as a logic variable selection problem and encode the selection results into a binary matrix $A\in\{0,1\}^{H\times(|\mathcal{X}|+M)}$. We then perform the continuous relaxations and relax $A$ to the selection probability $\tilde{A}\in[0,1]^{H\times(|\mathcal{X}|+M)}$.
  • (ii)
    Continuous Approximation of the Boolean Logic Features: We employ a logic-informed boolean feature approximation to ensure that the log-likelihood becomes a deterministic function of parameter $\tilde{A}$.
  • (iii)
    Differentiable Top-K Subset Sampling: We perform a differentiable top-K subset sampling with Gumbel noise injected to learn each rule. The learnable parameters are now becoming an unbounded weight matrix $W\in(R^{+})^{H\times(|\mathcal{X}|+M)}$. We leverage simulated annealing by tuning the temperature, which balances the approximation errors and issues caused by gradient vanishing in training.
In summary, we have performed the following series of approximation:
$\underbrace{\mathcal{F}}_{\text{rule set}}\to\underbrace{A}_{\text{binary}}\to\underbrace{\tilde{A}}_{\text{selection probability}}\to\underbrace{W}_{\text{unbounded weight}}$ (10)
and in the end we optimize $W$ end-to-end in a differentiable way. We provide a more detailed explanation as follows.
I: Continuous Relaxation of the Rule Learning
We first formulate the rule learning as a variable selection problem and introduce a binary matrix $A$ to encode the rule content (with columns indicating the logic variables and rows indicating the rules):
[…]
To accommodate for various rule lengths and meanwhile ensure that the rule length dones not exceed $K$, we append $M$ (a hyperparameter) dummy columns after $X_{|\mathcal{X}|}$, indicating empty-predicates to adjust for various rule lengths. We let $\sum_{j=1}^{|\mathcal{X}|+M}a_{hj}=K$, and whenever $a_{hj}=1$ for $j>|\mathcal{X}|$, it means that the rule length, determined by counting the valid predicates, will be less than $K$. A better illustration can be found in Fig. 2. Learning such $A$ involves challenging discrete optimization. Instead, we will relax the problem to learning the selection probability matrix $\tilde{A}$ with $0\leq\tilde{a}_{hj}\leq 1$.
Figure 2: Illustration of rule content. Top diagrams indicate $H$ learned rules given the bottom binary matrix $A$.
Given the selected predicates encoded in $A$, we can further identify their temporal relations. For each paired $X_{u}$ and $X_{v}$, we introduce a learnable parameter $\bm{\alpha}$ to denote their corresponding probabilities
$\bm{\alpha}=[\alpha_{b},\alpha_{e},\alpha_{a},\alpha_{n}]$
where $\alpha_{b}+\alpha_{e}+\alpha_{a}+\alpha_{n}=1$ and $0\leq\alpha_{b},\alpha_{e},\alpha_{a},\alpha_{n}\leq 1$. That is, $\bm{\alpha}$ belongs to a 3-d simplex. We soften the relation $r_{i}(X_{u},X_{v})$ by replacing it with
$\text{softmax}\left\{\alpha_{b}R_{b}\left(t_{u},t_{v}\right),\alpha_{e}R_{e}\left(t_{u},t_{v}\right),\alpha_{a}R_{a}\left(t_{u},t_{v}\right),\right.$
$\left.\alpha_{n}\left(1-\alpha_{b}R_{b}\left(t_{u},t_{v}\right)-\alpha_{e}R_{e}\left(t_{u},t_{v}\right)-\alpha_{a}R_{a}\left(t_{u},t_{v}\right)\right)\right\}.$ (12)
For each rule, we consider the temporal relations between all possible pairs of selected predicates. If a rule involves multiple combinations of two distinct predicates, we employ the softmin operation on these $r_{i}$ values. Using softmin helps prevent excessive value reduction resulting from multiple multiplications. The softmin value for a rule $h$ is denoted as ${{\mathcal{T}}}_{h},h=1,\dots,H$. The combination of softmax and softmin operations ensures that the probability learning process remains differentiable. We represent the selection probabilities for all temporal relations in a rule as $\bm{\alpha_{h}}$, and these learned values determine the types of temporal relations.
II: Continuous Approximation of the Boolean Logic Features
To learn the rule set in a differentiable way, in the M-step we propose an approximation to the boolean logic feature as shown in Eq. (3) and rewrite the intensity function as
$\lambda_{\text{soft}}^{*}(t_{i}\mid\bm{z}_{i})=b_{0}+\sum_{h=1,\dots,H}\gamma_{h}z_{ih}{{\mathcal{T}}}_{h}\mathcal{K}_{Lap}\left\{\sum_{j=1}^{|\mathcal{X}|}X_{j}(t_{j})\tilde{a}_{hj},K\right\}.$ (13)
We use Laplace kernel denoted as $\mathcal{K}_{Lap}$ with center $K$ to reparameterize the boolean feature and make it differentiable with $\tilde{A}$. As previously mentioned, ${{\mathcal{T}}}_{h}$ provides a softened approximation for temporal relations. This approach transforms the complete data log-likelihood into a deterministic function of $\tilde{A}$, enabling the gradient computation through automatic differentiation.
III: Differentiable Top-K Subset Sampling
Given the above description, now the rule learning process can be regarded as a top-K subset sampling process, which is governed by the selection probability $\tilde{A}$. We further use the reparameterization trick [18] for this top-K subset sampling.
Assume the top-K subset sampling can be parameterized by an unbounded weight matrix $W=[w_{1};w_{2};\dots;w_{H}]\in(R^{+})^{H\times(|\mathcal{X}|+M)}$, with each element non-negative. The top-$K$ relaxed subset sampling (using Gumbel noise) is summarized in Algorithm 1 (Appendix Section 1). By injecting the Gumbel noise, the introduced key $\hat{r}_{j}$ is differentiable with respect to the input weight $w_{h}$. We can use a top-$K$ relaxation based on successive applications of the softmax function [13] to obtain the relaxed top-K vector $\tilde{a}_{h}$.
To optimize sampling matrix $W$, we have
[…]
The parameter $W$ is updated with gradient descent.
With this reparameterization trick, the level of discreteness of $A$ can be controlled by the temperature parameter $\tau$. Thus, in our algorithm, an almost discrete matrix $A$ is used to select rules, and weight matrix $W$ is updated in every iteration to sample $A$. During the learning process, we employed cycled simulated annealing to fine-tune the temperature parameter $\tau$. The choice of $\tau$ balances the trade-off between approximation errors and gradient vanishing issues. Each time when we iterate all data, $\tau$ anneals from a large value to a small non-zero one.
5.2.2 Optimization wrt $\theta_{0}$
M-step: We optimize other model parameters by taking gradient of $\ell(\theta_{0},W)$ wrt $[b_{0},[r_{h}]_{h\in[1,\dots,H]},[\bm{\alpha_{h}}]_{h\in[1,\dots,H]},\bm{\pi}]$, and set these gradients to 0. We need to make sure that these model parameters are non-negative. This can be guaranteed by projected gradient descent or using the change-of-variable trick (e.g., parameterize $b_{0}$ as $\exp(b_{0})$).
Specifically, solving $\bm{\pi}$ in the M-step has a closed-form solution. Given the constraints $\sum_{h=0}^{H}\pi_{h}=1$ we have
[…]
6 Experiments
6.1 Synthetic Data Experiments
We assess our model’s learning accuracy in terms of rule discovery, rule weight estimation, and rule prior distribution learning across various scenarios.
I: Experiment Setup
In our empirical analysis, we systematically explore various aspects to evaluate the scalability and performance of our method: (i) Diverse Ground Truth Rules: We vary the number of ground truth rules, ranging from 1 to 4, to examine the model’s performance across different complexity levels. (ii) Variation in Predicate Set: We consider five scenarios with different size of predicate set, including 10, 15, 20, 25, and 30, to assess the scalability of our method. A larger logic variable set represents a more challenging problem. (iii) Variation in Data Size: We conduct experiments with varying data sizes, including sequences of 5000, 10,000, 20,000, 30,000, and 40,000 instances, to investigate the EM algorithm’s scalability under different data scales.
It is important to note that only a small subset of the predicates is excluded in the ground truth rules. In other words, the majority of the candidate logic variables are dummy variables. We aim to demonstrate that our method excels in pinpointing the most critical logic variables during the rule learning process. We present the primary results based on experiments with 20,000 sequences for each scenario. For additional details and comprehensive findings, please refer to the Appendix (Section 2).
Figure 3: Our proposed model’s performance is evaluated across all five scenarios for group-2, with two ground truth rules. We evaluate our model’s performance in terms of rule discovery, rule weight learning, and rule prior distribution learning. The color “blue” indicates ground truth rules, weights, and prior distributions, whereas the colors “red” and “yellow” indicate the learning results.
Figure 4: Jaccard similarity score and MAE of rule weights and rule preferential probabilities for all 4 groups. X-axis indicates the predicate library size and Y-axis indicates the value of Jaccard similarity and MAE.
II: Accuracy
We evaluate the accuracy of our method in terms of rule set discovery and point process model parameter learning. In particular, we assess its accuracy in increasingly challenging scenarios by gradually enhancing the number of dummy logic variables in the predicate set.
In Fig. 3, we present the results for group-2, while the complete results are available in the Appendix (Section 2). The top row of plots employs Venn diagrams to visually depict the agreement between the actual rule set and the learned rule set, quantified through the Jaccard similarity score (the intersection area divided by the union area). Our proposed model consistently identifies almost all the ground truth rules across all groups.
The middle row diagrams compare the true rule weights with the learned rule weights, and the bottom row diagrams compare the true rule prior distributions with the learned rule distributions. Once the ground truth rules are accurately learned, the learned prior rule distributions closely align with the true values, resulting in low mean absolute errors (MAEs).
In addition, we also compared our method with several classic unsupervised rule mining methods, including Apriori [1], NEclatcloesed [2], CM-SPADE [6] and VGEN [7]. These baseline methods identify rules using the frequency thresholds, resulting in coarse rule discovery. From the experiment results, these approaches lead to the extraction of a vast number of noisy and erroneous rules, while the few accurate rules tend to remain hidden within this extensive rule set. A complete results and discussion can be found in Appendix (Section 4 and 6).
III: Scalability
Fig. 4 illustrates the Jaccard similarity scores and MAE for all groups as the number of redundant predicates increases. Across all four groups, we observe a decrease in Jaccard similarity score and a slight increase in MAE with the growing number of predicates. However, it’s important to note that our model maintains stable and reliable when the size of the predicate set is appropriate (i.e., around 20-30) and an adequate event sequences are available.
Furthermore, our approach exhibits promising results in identifying the causal rules for each event, inferred by the posterior of the latent rule index. Tab. 1 presents the accuracy of the posterior probabilities for each rule. Notably, in scenarios where almost all ground truth rules were successfully learned, the posterior accuracy is exceptionally high. However, in more challenging tasks, such as those in Group-3 and Group-4, where a larger number of ground truth rules are involved, the accuracy of the rule index posterior experiences a slight decrease.
In fact, our proposed model didn’t deviate significantly in the case of erroneously uncovered rules. Fig. 5 provides an example of a learned wrong rule from Group-3, Case-4. In this instance, the discovered inaccurate rule includes only one additional redundant predicate. Similar patterns are observed in other cases as well.
Figure 5: One example of erroneously uncovered rule for Group-3 (3 ground truth rules) Case-4 (23 to-be-searched predicates). Ground truth rule: $Y\leftarrow x_{4}\land x_{5}\land(x_{4}\,\text{Before}\,x_{5})$. Learned rule: $Y\leftarrow x_{1}\land x_{4}\land x_{5}\land(\ x_{4}\,\text{Before}\ x_{5})$. We see that only one more predicate was wrongly excluded.
Case Group Group-1 Group-2 Group-3 Group-4
1 (10preds) 0.8036 0.7794 0.2818 0.2658
2 (15preds) 0.7962 0.7831 0.4180 0.5155
3 (20preds) 0.7979 0.7717 0.4224 0.4959
4 (25preds) 0.7986 0.6821 0.2802 0.5557
5 (30preds) 0.8039 0.6789 0.2675 0.4960
Table 1: Rule index inference accuracy on synthetic data for all 4 groups.
IV: Event Prediction
We compared our model with several state-of-the-art baselines in terms of event predictions, using MAE as the evaluation metric for event time prediction. A complete descriptions of these baseline methods can be found in Appendix Section 4. As shown in Tab. 2, our model outperforms all baselines, particularly in complex scenarios like case-5 with 30 predicates across all groups. In other cases, the MAE varies depending on rule learning outcomes, but it consistently matches or surpasses the baselines.
Method Group Group-1 Group-2 Group-3 Group-4
THP[23] 2.2155 2.0089 2.2364 2.4121
RMTPP[4] 2.2354 2.1843 2.3147 2.4523
ERPP[17] 2.3532 2.0175 2.3564 2.4264
GCH[19] 2.5233 2.3754 2.4655 2.6427
LG-NPP[22] 2.4145 2.4732 2.6564 2.5612
GM-NLF[5] 2.3147 2.3622 2.4522 2.4362
TELLER[10] 2.2281 2.4812 2.5121 2.4030
CLNN[20] 2.3858 2.5165 2.4972 2.5096
OURS* 2.2075 1.8586 2.1957 2.3559
Table 2: Event time prediction MAE on synthetic data for case-5 (30 preds) of all 4 groups.
6.2 Healthcare Data Experiments
MIMIC-IV [^1] https://mimic.mit.edu/ is a dataset of electronic health records for patients who have been admitted to the intensive care unit (ICU) [8]. In our study, we focused on 4074 patients diagnosed with sepsis [15], as sepsis is a leading cause of mortality in the ICU, particularly when it progresses to septic shock. Septic shocks are critical medical emergencies, and timely recognition and treatment are crucial for improving survival rates. The objective of this experiment is to identify logic rules associated with septic shocks for both the general population and specific patients. These rules could serve as potential early alarms when abnormal indicators are detected, aiding in timely intervention.
Weight Rule
0.566
Rule 1: LowUrine $\leftarrow$
Arterial Blood
Pressure Diastolic
0.467
Rule 2: LowUrine $\leftarrow$
Inoized Calcium
0.527
Rule 3: LowUrine $\leftarrow$
Arterial CO2 Pressure
0.466
Rule 4: LowUrine $\leftarrow$
Venous 02 Pressure
0.464
Rule 5: LowUrine $\leftarrow$
Respiratory Rate $\wedge$ Hemoglobin
0.664
Rule 6: LowUrine $\leftarrow$ Heart Rate $\wedge$ BUN $\wedge$
WBC $\wedge$ (Heart Rate Before BUN)
$\wedge$ (BUN Before WBC)
Table 3: Some examples of our discovered temporal logic rules in MIMIC-IV.
Routine Vital Signs and Lab values
A total of 29 variables associated with sepsis were extracted, including routine vital signs and laboratory values, as suggested and utilized in [9]. We recorded the time points at which these variables first exhibited abnormal values within the 48-hour period preceding the onset of abnormal urine output (i.e., our target event).
Discovered Logic Rules
Real-time urine output was utilized as the health state indicator, as low urine output directly indicates a deteriorate circulatory system and serves as a warning sign for septic shock. Due to the frequent fluctuations in urine output within the ICU setting, only instances where urine output becomes abnormal after maintaining a normal level for at least 48 hours were considered valid target events (i.e., meaningful to predict and explain).
In Tab. 3, we present a portion of the discovered explanatory logic rules and their corresponding learned weights based on our method. These displayed rules have been justified by doctors.
Rule 1 highlights the significance of abnormal arterial blood pressure diastolic as a crucial indicator for septic shock, with a substantial rule weight of 0.566. Therefore, patients should closely pay high attention to their arterial blood pressure diastolic and promptly seek clinical support if it deviates from the normal range.
Rule 2, on the other hand, noted by doctors, may be less relevant as fluctuations in calcium homeostasis can occur in specific patients during sepsis. Rule 3 and Rule 4 highlight the individual significance of abnormal arterial $CO_{2}$ pressure and venous $O_{2}$ pressure as additional early indicators of a compromised circulatory system. These rules bear relatively significant weights of 0.527 and 0.466, respectively.
Rule 5 and Rule 6 demonstrate the combined effects on low urine output. Rule 5, with a weight of 0.464, suggests that when the respiratory rate of patients exceeds 20 breaths per minute and there is a sharp decline in hemoglobin levels, they should be highly vigilant for the possibility of septic shock. Furthermore, Rule 6, which has the highest weight of 0.664, indicates that if the heart rate of patients surpasses 90 beats per minute, blood urea nitrogen (BUN) and white blood cell count (WBC) fall outside the normal ranges, and all three signals occur successively, immediate emergency medical treatment should be sought.
Our methodology incorporated invaluable input from medical experts who thoroughly reviewed our discovered rule results. Their feedback consistently validated the soundness and relevance of our findings. Additionally, we bolstered our discovered rules by referencing and citing supporting evidence from reputable medical sources, with discussions found in Appendix Section 8. We also compared our method with the post-hoc method [3] in terms of interpretability, with the corresponding results in Appendix Section 5.
Figure 6: Rule identification results for patient 14823401. X-axis: time ranges from 0 to 48 h. Y-axis: feature id (29 in total). Low urine output occurred at 48h. Deep blue blocks record the occurrence time of a routine vital sign or abnormal lab value. Learned rule with largest posterior probability (colored in purple): LowUrine $\leftarrow$ Heart Rate $\land$ BUN $\land$ WBC, Heart Rate Before BUN, BUN Before WBC. Other rules with relatively large posterior probability are colored in pink and off white. The darker the color, the greater the posterior probability.
Logic Rule Identification for Each Event
Our method can identify the most likely causal rule specific to individual patients. For instance, let’s consider patient with ID:14823401 as an example and the results are visualized in Fig. 6. Based on the analysis, the rule with the highest posterior probability that the patient is likely to adhere to is Rule 6: LowUrine $\leftarrow$ Heart Rate $\land$ BUN $\land$ WBC. The identified rule will help with the individual healthcare delivery and treatment design.
Event Prediction
We used the same baselines as in the synthetic data experiments, with the MAE serving as the evaluation metric for predicting the occurrence time of “Low Urine” events. Our model’s performance is presented in Tab. 4, where it outperforms all the baselines.
Method LowUrine
THP 2.4234
RMTPP 2.4643
ERPP 2.6122
GCH 2.5367
LG-NPP 2.5672
GM-NLF 2.6925
TELLER 2.4401
CLNN 2.4371
OURS* 2.3675
Table 4: Event time prediction MAE on healthcare data.
7 Conclusion
In this paper, we leverage the temporal point process models to effectively discover latent logic rules to predict and explain abnormal events. We design an EM algorithm that can jointly learn model parameters, discover rule set, and infer the most likely rule factor for each event. We especially can learn the rule set in a differentiable way. Our method has exhibited strong performance, as demonstrated through experiments with synthetic and real healthcare datasets.
Checklist
  1. 1.
    For all models and algorithms presented, check if you include:
    1. (a)
      A clear description of the mathematical setting, assumptions, algorithm, and/or model. [Yes]
    2. (b)
      An analysis of the properties and complexity (time, space, sample size) of any algorithm. [Yes]
    3. (c)
      (Optional) Anonymized source code, with specification of all dependencies, including external libraries. [Yes]
  2. 2.
    For any theoretical claim, check if you include:
    1. (a)
      Statements of the full set of assumptions of all theoretical results. [Yes]
    2. (b)
      Complete proofs of all theoretical results. [Yes]
    3. (c)
      Clear explanations of any assumptions. [Yes]
  3. 3.
    For all figures and tables that present empirical results, check if you include:
    1. (a)
      The code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL). [Yes]
    2. (b)
      All the training details (e.g., data splits, hyperparameters, how they were chosen). [No]
    3. (c)
      A clear definition of the specific measure or statistics and error bars (e.g., with respect to the random seed after running experiments multiple times). [No]
    4. (d)
      A description of the computing infrastructure used. (e.g., type of GPUs, internal cluster, or cloud provider). [Yes]
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets, check if you include:
    1. (a)
      Citations of the creator If your work uses existing assets. [Not Applicable]
    2. (b)
      The license information of the assets, if applicable. [Not Applicable]
    3. (c)
      New assets either in the supplemental material or as a URL, if applicable. [Not Applicable]
    4. (d)
      Information about consent from data providers/curators. [Not Applicable]
    5. (e)
      Discussion of sensible content if applicable, e.g., personally identifiable information or offensive content. [Not Applicable]
  5. 5.
    If you used crowdsourcing or conducted research with human subjects, check if you include:
    1. (a)
      The full text of instructions given to participants and screenshots. [Not Applicable]
    2. (b)
      Descriptions of potential participant risks, with links to Institutional Review Board (IRB) approvals if applicable. [Not Applicable]
    3. (c)
      The estimated hourly wage paid to participants and the total amount spent on participant compensation. [Not Applicable]